OpenStudy (lukecrayonz):

How do i describe the shape of the graph of each cubic function by determining the end behavior and number of turning points? for ex. y=3x^3-x-3

OpenStudy (mathmate):

If the leading coefficient (coef. of x^3) is positive, the graph goes from on the left -inf to +inf. If it is negative, it goes from +inf (on the left) to -inf. Check f'(x)=0. If there are two real and distinct solutions, there will be a maximum and a minimum. If the two solutions are coincident, then there is a turning point (neither max nor min). If there is no solution, the the graph goes monotonic inc. or dec. You can also describe the zeroes by solving for them.

OpenStudy (anonymous):

dy/dx and dy²/d²x

OpenStudy (anonymous):

dy/dx = 9x²-1

OpenStudy (anonymous):

when dy/dx=0, u will find its turning point 9x²-1=0 x²=1/9 x=(+ or -) 1/3 tat mean the truing point of the graph has x of -1/3 and -1/3. to find the y of the turning point, simply input the value of x into the original function in ur case: when x=-1/3 y=3(-1/3)^3 +1/3 - 3 when x =1/3 y=3(1/3)^3 -1/3 - 3